OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (texaschic101):

a + s = 150 --> a = 150 - s 45a + 28.5s = 4770 now sub 150 - s in for a in the 2nd equation 45(150 - s) + 28.5s = 4770 6750 - 45s + 28.5s = 4770 - 16.5s = 4770 - 6750 - 16.5s = - 1980 s = - 1980/- 16.5 s = 120 Now sub 120 in for s in the 1st equation a + s = 150 a + 120 = 150 a = 150 - 120 a = 30 check.... 45a + 28.5s = 4770 45(30) + 28.5(120) = 4770 1350 + 3420 = 4770 4770 = 4770 (correct) ANSWER : There were 30 adults and 120 students.

OpenStudy (anonymous):

thanks i clearly see it better now

OpenStudy (texaschic101):

your welcome :)